HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 71     Entry time: Wed Apr 2 08:47:27 2025
Author: Michael, Alex 
Category: General 
Subject: The beam is back 
The beam is back. But, the experts have found a problem with one of the modules of the SIS kicker magnet. And, currently they have disabled this module. This fixed the problem. However, since the kicker system is not running with full power, it is not clear by how much this can affect the extracted beam intensity for the ESR. Therefore,  the ESR experts have been called and they will look at the injected beam intensity in the ring, in order to know if its still the same as before disabling the module. 
The experts will discuss at the upcoming meeting how to proceed with the damaged module, i.e. when to repair it.
In addition, the FRS started to take the beam now (again) which means that we will be getting less pulses from SIS as compared to the previous night.
ELOG V3.1.5-fc6679b